ok i am in Algebra A and i am doing weighted averages. here is the problem: An Airplane flies 1000 miles due east in 2 hours and 1000 miles due south in 3 hours.what is the average speed of the airplane.
Answer by checkley71(8403)

THE TOTAL MILEAGE IS 2,000 MILES THAT TAKES 2+3=5 HOURS.
THUS WE HAVE THE FOLLOWING EQUATION TO FIND THE AVERAGE SPEED
(1000+1000)/(2+3)=2000/5=400 MPH IS THE AVERAGE SPEED.